# Backstepping for Partial Differential Equations

### v2

@@ 4.4 Nonlinear PDEs
However, it inherits fundamental limitations of linearization. Stability properties are, at best, valid only for small perturbations around the linearization point, and in some instances, the linearized system can be exponentially stable while the nonlinear PDE system lacks even local stability. In certain cases, a linearizing transformation may exist that allows for the conversion of the nonlinear PDE into a linear one. This approach, showcased in [189, 188] for the Burgers equation and in [53] for a Hamilton-Jacobi nonlinear PDE, relies on finding a suitable linearizing change of variables, which in this case is the Cole-Hopf transformation [135, p. 206]. The resulting linear system can then be stabilized using standard backstepping methods, yielding a controller that is nonlinear in the original state. The success of this approach hinges on the existence of such a transformation, which is not guaranteed for all nonlinear PDEs and is typically not global. For systems involving a nonlinear ODE coupled with a PDE, backstepping can sometimes be employed to compensate for the PDE dynamics and achieve stabilization. Several examples are presented in the works [50, 48, 52], where the authors address the problem of compensating wave or delay dynamics for nonlinear ODE systems. The design involves a predictor-based feedback law, compensating for the dynamics introduced by the different PDEs. Interestingly, in some cases, a linear feedback law may suffice to stabilize a nonlinear PDE. In [165], the authors present a global stabilization result for a class of nonlinear reaction-diffusion PDEs using boundary feedback. The key insight is that for certain nonlinearities, a simple linear feedback law can induce a “damping” effect when the norm of the state becomes large, leading to global exponential stability in the norm. However, the most general and ambitious approach for nonlinear PDE control is the one based on spatial Volterra series444A Volterra series is a functional series representation of a nonlinear system, generalizing the concept of a single Volterra integral. While a single Volterra integral captures the output of a system as a weighted sum of input values, a Volterra series extends this notion by incorporating multiple integrals of increasing order, capturing higher-order nonlinear interactions between input values.. The works [268, 269] introduce a framework that employs spatial Volterra series nonlinear operators both in the state transformation and the feedback law, generalizing the concept of feedback linearization to infinite dimensions. This approach allows for the stabilization of a broad class of nonlinear parabolic PDEs, but its complexity and the inherent challenges in proving the global invertibility of the transformation (in general, only local invertibility can be proved) have limited its application to specific cases.
@@ 4.5 Time-Varying PDEs and Prescribed/Fixed-Time Stabilization
The extension of backstepping to time-varying PDEs, where the coefficients in the PDE depend on time, introduces significant challenges compared to the time-invariant case beyond a few cases where explicit solutions are available [248]. The primary obstacle stems from the complex interplay between the time-varying coefficients and the spatial dynamics of the PDE. For parabolic PDEs, [279, 166] demonstrate that the backstepping method can be applied to systems with time-varying coefficients, but the stability analysis requires imposing restrictions on the regularity of these coefficients. Specifically, the time-varying parameters are typically assumed to belong to certain Gevrey classes, which guarantees the existence of solutions for the kernel equations with specific smoothness properties. This reliance on Gevrey regularity stems from the inherent smoothing properties of parabolic PDEs, where the solution at any time instant is influenced by the entire past history of the system. An implementation to solve the resulting kernel equations is given in [160]. In contrast, for hyperbolic PDEs, the situation is less restrictive, as highlighted in [118, 14, 94, 89]. The finite speed of propagation inherent to hyperbolic dynamics allows for a more localized influence of the time-varying coefficients. As a result, backstepping designs for time-varying hyperbolic PDEs can be developed without relying on Gevrey regularity, as the solution at a given point in spacetime is affected only by a limited portion of the time-varying parameter history. Recently, alternative gain-scheduling approaches have been proposed in [161, 35]. Instead of handling time-varying backstepping kernels that account for both temporal and spatial variations in the coefficients, these approaches consider only the spatial variations of the coupling coefficients. This method remains valid as long as the coupling coefficients are sampled over time, thereby reducing the kernel PDEs to involve only space-dependent coefficients between successive sampling or triggering instants. These methods do not require solving time-varying backstepping kernel equations in real-time resulting in reduced computational burden and broader applicability.
@@
The goal of achieving stabilization within a prescribed time, independent of initial conditions, further complicates the design. This is where the concepts of fixed-time and prescribed-time stabilization become relevant. Finite-time stability refers to systems that reach equilibrium within a finite time, but this time depends on the initial conditions. Fixed-time stability, a stronger notion, guarantees convergence to the equilibrium within a time that is uniformly bounded with respect to the initial conditions. In prescribed-time stability, the settling time can be explicitly prescribed a priori, independently of both initial conditions and system parameters. For parabolic PDEs, achieving finite-time or fixed-time stability typically requires the use of time-varying feedback gains. In [92], this is achieved for space-varying reaction-diffusion systems by constructing a sequence of backstepping kernels mapping the plant into a sequence of target systems with successively increased damping coefficients, which allows to design periodic time-varying feedback laws. Alternatively, in [133], a single backstepping transformation with a time-varying kernel and a single target system with a time-varying coefficient is considered, ensuring convergence within a predetermined time, but the approach only covers constant-coefficient reaction-diffusion systems. Similarly, [134, 132] focuses on prescribed-time stabilization and estimation of LTI systems with input delay, employing a time-varying predictor feedback based on a transport PDE representation of the delay and time-varying backstepping kernels. These approaches rely on selecting a time-varying target system that exhibits the desired finite-time convergence property. For parabolic PDEs, this is essential because these systems do not naturally exhibit finite-time convergence to the origin. In contrast, hyperbolic equations can achieve finite-time convergence to zero, but with a minimum time determined by the propagation speed of the characteristics. This distinction highlights the fundamental difference between the temporal behavior of these two classes of PDEs. In a related context, [255] presents prescribed-time observers and controllers for the linearized Schrödinger equation using time-varying gains, also based on backstepping. The observer gain diverges as time approaches the prescribed terminal time, but this growth is carefully managed to guarantee prescribed–time convergence of the estimate. The authors also demonstrate prescribed–time output regulation by combining the estimation and control results.
@@ 4.6 Moving Boundary Domains
The control and estimation of PDEs with time-varying spatial domains, known as moving boundary problems, present considerable challenges beyond those encountered in fixed-domain scenarios. The moving boundaries introduce time-dependent spatial domains and boundary conditions, leading to a more complex mathematical framework and necessitating specialized control and observer designs. Backstepping has shown remarkable flexibility in adapting to these challenges, with several contributions showcasing its efficacy in tackling moving boundary problems. One of the earliest works in this domain is [158], an output feedback control design for moving boundary parabolic PDEs is proposed using backstepping. The authors introduce a new coordinate transformation that incorporates both the state of the system and the moving boundary, leading to a fixed-domain target system but a time-varying PDE. A backstepping observer is designed to estimate the state using boundary measurements, and the estimated state is then used in a feedback law that compensates for the moving boundary effects. The problem of boundary stabilization of one-dimensional cross-diffusion systems in a moving domain is studied in [73]. In [168] the authors address the stabilization of the one-phase Stefan problem, a classic moving boundary problem describing melting or solidification processes. They employ a backstepping transformation that explicitly accounts for the moving boundary, mapping the Stefan problem into a stable target system with a fixed boundary. This approach leads to a boundary controller that effectively regulates the interface position and ensures exponential stability in a suitable norm. Building upon this work, [171] extends the backstepping design to the broader class of two-phase Stefan problems. Recently, [257] explores adaptive control in moving boundary problems for the wave equation, by rewriting the problem into an LTI form with time-varying input delay and unknown parameters to which adaptive backstepping designs can be applied. See also Section 5.2 for some applications of these designs.
@@ 4.7 Fredholm operators and non-strict feedback plants
As mentioned in the primer, in Section 2, while the original backstepping method was developed using Volterra integral transformations, see (8), in some cases this does not give enough degrees of freedom to find an adequate target system and well-posed kernel equation for the control design. The first specific example where a different type of transform was looked for is likely [245], to assign an arbitrary decay rate for an Euler-Bernoulli beam equation. This early example builds a control law based on a previous design for the Schrödinger equation, which uses a standard Volterra transformation, and then adapts it to obtain stabilization to zero of the Euler-Bernoulli beam equation instead of to a constant. Unlike later backstepping designs, this design depends on the system dynamics when defining the properties of the transforms, these are not seen as operators mapping from generic function spaces (e.g. to ), but relating solutions to two specific differential equations. This can be seen by the fact that the time derivative of the variable appears on one side of the transformations. In this particular case, the Fredholm transformation linking the two sets of solutions is found after imposing a form on the control action in order to obtain explicit expressions for the closed-loop system. The next examples found in the literature would be[45] with a wave PDE in the actuation path of an ODE and [47] in the case of distributed input delays in an ODE, where the direct backstepping transformation is written in a more general form. Nevertheless, this transformation, due to the particular structure of the system under consideration, has good invertibility properties. In particular, the inverse transform only contains Volterra integrals, which seems to indicate that the transformation is not completely Fredholm. Another particular structure, this time concerning parabolic equations with integral terms, was presented in [149] for a spatially non-causal reaction-diffusion equation. In this case, the particular form of the integral kernel in the non-local term is exploited to obtain an explicit solution for the Fredholm transformation and its invertibility. After these first examples, two different early approaches emerged to tackle more general Fredholm transformations, a contraction approach in [67] and [68] that required little structure from the coefficients yet only worked for “small” coefficients; and a more technically involved approach tailored to self-adjoint and skew-adjoint operators in the case of a Korteweg-de Vries equation in [90] and a Kuramoto-Sivashinsky equation in [91], respectively. In these last two papers, the existence and invertibility of the transform were related to controllability conditions. Many of the more recent results follow more closely a later approach, presented in [87], which linked the existence and invertibility of the Fredholm transform to a controllability condition in the case of first-order differential equations, yet required less specific conditions on the operators studied.
@@ 4.8 Sampled-data and Event-triggered systems
As discussed in section 4.1, one of the strengths of the backstepping method has been its capacity to simplify convoluted structures in PDE systems and allow for an equivalent stability analysis based on a simpler target system. This advantage has been exploited not only in adaptive and parameter estimation methods, but also in sampled-data and Event-triggered stabilization and estimation schemes, which generally exploit the easier structure of the target system in order to construct Lyapunov functions for the closed-loop system (or to estimate the rate of decrease of some norm of the original system).
@@
Event-based control results include linear hyperbolic systems in [130], as well as an observer-based version in [128]. Another related system can be found in [129], where traffic flow on cascaded roads is considered under event-triggered output-feedback. As for parabolic PDEs, existing results include [131] for reaction-diffusion PDEs with constant coefficients, as well as [229] for an observer-based control and [230] including also sampled-data. For interconnected PDE-ODE systems, some results also exist, such as [204] in the hyperbolic case, and [285] in the parabolic case.
@@ 5.1 Traffic Control
Freeway traffic modeling and control have been intensively investigated over the past decades. Backstepping has proven to be an effective method for designing boundary controllers that address stop-and-go oscillations in congested traffic, often modeled by the nonlinear Aw-Rascle-Zhang (ARZ) model [301, 299]. This model, consisting of second-order hyperbolic PDEs, accurately captures the complex dynamics of traffic density and velocity. Early backstepping designs focused on stabilizing traffic flow on a single freeway segment, considering both upstream and downstream control problems using ramp metering [301]. They employed backstepping transformations to map the original system to a target system with desirable stability properties, yielding full-state feedback control laws and collocated/anti-collocated observers. These designs achieved exponential stability and finite-time convergence to equilibrium, validated through simulations. Later works expanded the methodology to address the more challenging problem of simultaneous downstream and upstream stabilization on cascaded freeway segments connected by a junction, considering varying road conditions [299]. This scenario requires a more sophisticated control approach due to the interconnected nature of the system and distinct traffic scenarios on different segments. Backstepping-based output-feedback controllers were designed to suppress stop-and-go oscillations, leveraging boundary measurements of flow rate and velocity at the junction or the outlet of the segment. These designs ensured robust stabilization of the under-actuated network of two hyperbolic PDE systems, effectively mitigating congestion on cascaded roads. In [147, 148] these works were further extended the application of backstepping to the control of multi-class traffic flow, incorporating the influence of disturbances and bottlenecks they designed an optimal observer-based output feedback controller for a heterogeneous traffic flow model with disturbances at the inlet of a considered road segment. Leveraging the backstepping approach, the proposed controller guaranteed the integral input-to-state stability of the closed-loop system and effectively addressed traffic breakdown caused by a combination of high traffic demand, bottlenecks, and driver disturbances. The controller design involved constructing an observer that estimates the states of the linearized ARZ traffic flow model using only boundary measurements. In [146], the backstepping method was employed to design an observer for a heterogeneous quasilinear traffic flow system, explicitly addressing the impact of disturbances at the inlet of the road section. This work demonstrated the applicability of backstepping for state observation in the presence of perturbations, ensuring accurate estimation of the traffic state despite uncertainties and nonlinearities. Backstepping has also been applied to address the control of traffic flow with multiple vehicle classes, each characterized by its own size and driver behavior [69]. This requires a multi-class traffic model, typically represented by a set of coupled hyperbolic PDEs. By designing an output-feedback controller based on backstepping and an anti-collocated observer, researchers achieved finite-time damping of stop-and-go waves by measuring the velocities and densities of different vehicle classes at the inlet of the considered track section. In another work, a backstepping observer was developed and validated with real freeway data to estimate congested freeway traffic states based on the ARZ model [300]. This observer design only requires boundary measurements of flow and velocity, making it suitable for practical applications with limited sensor coverage.

### v3

@@ 4.4 Nonlinear PDEs
However, it inherits fundamental limitations of linearization. Stability properties are, at best, valid only for small perturbations around the linearization point, and in some instances, the linearized system can be exponentially stable while the nonlinear PDE system lacks even local stability.
In certain cases, a linearizing transformation may exist that allows for the conversion of the nonlinear PDE into a linear one. This approach, showcased in [190, 189] for the Burgers equation and in [53] for a Hamilton-Jacobi nonlinear PDE, relies on finding a suitable linearizing change of variables, which in this case is the Cole-Hopf transformation [136, p. 206]. The resulting linear system can then be stabilized using standard backstepping methods, yielding a controller that is nonlinear in the original state. The success of this approach hinges on the existence of such a transformation, which is not guaranteed for all nonlinear PDEs and is typically not global. For systems involving a nonlinear ODE coupled with a PDE, backstepping can sometimes be employed to compensate for the PDE dynamics and achieve stabilization. Several examples are presented in the works [50, 48, 52], where the authors address the problem of compensating wave or delay dynamics for nonlinear ODE systems. The design involves a predictor-based feedback law, compensating for the dynamics introduced by the different PDEs. Interestingly, in some cases, a linear feedback law may suffice to stabilize a nonlinear PDE. In [166], the authors present a global stabilization result for a class of nonlinear reaction-diffusion PDEs using boundary feedback. The key insight is that for certain nonlinearities, a simple linear feedback law can induce a “damping” effect when the norm of the state becomes large, leading to global exponential stability in the norm. However, the most general and ambitious approach for nonlinear PDE control is the one based on spatial Volterra series444A Volterra series is a functional series representation of a nonlinear system, generalizing the concept of a single Volterra integral. While a single Volterra integral captures the output of a system as a weighted sum of input values, a Volterra series extends this notion by incorporating multiple integrals of increasing order, capturing higher-order nonlinear interactions between input values.. The works [271, 272] introduce a framework that employs spatial Volterra series nonlinear operators both in the state transformation and the feedback law, generalizing the concept of feedback linearization to infinite dimensions. This approach allows for the stabilization of a broad class of nonlinear parabolic PDEs, but its complexity and the inherent challenges in proving the global invertibility of the transformation (in general, only local invertibility can be proved) have limited its application to specific cases.
@@ 4.5 Time-Varying PDEs and Prescribed/Fixed-Time Stabilization
The extension of backstepping to time-varying PDEs, i.e. where the coefficients in the PDE depend on time, introduces significant challenges compared to the time-invariant case. Beyond a few cases where explicit solutions are available [250], the primary obstacle stems from the complex interplay between the time-varying coefficients and the spatial dynamics of the PDE. For parabolic PDEs, [284, 167] demonstrate that the backstepping method can be applied to systems with time-varying coefficients, but the stability analysis requires imposing restrictions on the regularity of these coefficients. Specifically, the time-varying parameters are typically assumed to belong to certain Gevrey classes, which guarantees the existence of solutions for the kernel equations with specific smoothness properties. This reliance on Gevrey regularity stems from the inherent smoothing properties of parabolic PDEs, where the solution at any time instant is influenced by the entire past history of the system. An implementation to solve the resulting kernel equations is given in [161]. In contrast, for hyperbolic PDEs, the situation is less restrictive, as highlighted in [119, 14, 95, 90]. The finite speed of propagation inherent to hyperbolic dynamics allows for a more localized influence of the time-varying coefficients. As a result, backstepping designs for time-varying hyperbolic PDEs can be developed without relying on Gevrey regularity, as the solution at a given point in spacetime is affected only by a limited portion of the time-varying parameter history. Recently, alternative gain-scheduling approaches have been proposed in [162, 35]. Instead of handling time-varying backstepping kernels that account for both temporal and spatial variations in the coefficients, these approaches consider only the spatial variations of the coupling coefficients. This method remains valid as long as the coupling coefficients are sampled over time, thereby reducing the kernel PDEs to involve only space-dependent coefficients between successive sampling or triggering instants. These methods do not require solving time-varying backstepping kernel equations in real-time resulting in reduced computational burden and broader applicability.
@@
The goal of achieving stabilization within a prescribed time, independent of initial conditions, further complicates the design. This is where the concepts of fixed-time and prescribed-time stabilization become relevant. Finite-time stability refers to systems that reach equilibrium within a finite time, but this time depends on the initial conditions. Fixed-time stability, a stronger notion, guarantees convergence to the equilibrium within a time that is uniformly bounded with respect to the initial conditions. In prescribed-time stability, the settling time can be explicitly prescribed a priori, independently of both initial conditions and system parameters. For parabolic PDEs, achieving finite-time or fixed-time stability typically requires the use of time-varying feedback gains. In [93], this is achieved for space-varying reaction-diffusion systems by constructing a sequence of backstepping kernels mapping the plant into a sequence of target systems with successively increased damping coefficients, which allows to designing periodic time-varying feedback laws. Alternatively, in [134], a single backstepping transformation with a time-varying kernel and a single target system with a time-varying coefficient is considered, ensuring convergence within a predetermined time, but the approach only covers constant-coefficient reaction-diffusion systems. Similarly, [135, 133] focuses on prescribed-time stabilization and estimation of LTI systems with input delay, employing a time-varying predictor feedback based on a transport PDE representation of the delay and time-varying backstepping kernels. These approaches rely on selecting a time-varying target system that exhibits the desired finite-time convergence property. For parabolic PDEs, this is essential because these systems do not naturally exhibit finite-time convergence to the origin. In contrast, hyperbolic equations can achieve finite-time convergence to zero, but with a minimum time determined by the propagation speed of the characteristics. This distinction highlights the fundamental difference between the temporal behavior of these two classes of PDEs. In a related context, [257] presents prescribed-time observers and controllers for the linearized Schrödinger equation using time-varying gains, also based on backstepping. The observer gain diverges as time approaches the prescribed terminal time, but this growth is carefully managed to guarantee prescribed–time convergence of the estimate. The authors also demonstrate prescribed–time output regulation by combining the estimation and control results.
@@ 4.6 Moving Boundary Domains
The control and estimation of PDEs with time-varying spatial domains, known as moving boundary problems, present considerable challenges beyond those encountered in fixed-domain scenarios. The moving boundaries introduce time-dependent spatial domains and boundary conditions, leading to a more complex mathematical framework and necessitating specialized control and observer designs. Backstepping has shown remarkable flexibility in adapting to these challenges, with several contributions showcasing its efficacy in tackling moving boundary problems. One of the earliest works in this domain is [159], where an output feedback control design for moving boundary parabolic PDEs is proposed using backstepping. The authors introduce a new coordinate transformation that incorporates both the state of the system and the moving boundary, leading to a fixed-domain target system but a time-varying PDE. A backstepping observer is designed to estimate the state using boundary measurements, and the estimated state is then used in a feedback law that compensates for the moving boundary effects. The problem of boundary stabilization of one-dimensional cross-diffusion systems in a moving domain is studied in [74]. In [169] the authors address the stabilization of the one-phase Stefan problem, a classic moving boundary problem describing melting or solidification processes. They employ a backstepping transformation that explicitly accounts for the moving boundary, mapping the Stefan problem into a stable target system with a fixed boundary. This approach leads to a boundary controller that effectively regulates the interface position and ensures exponential stability in a suitable norm. Building upon this work, [172] extends the backstepping design to the broader class of two-phase Stefan problems. Recently, [259] explores adaptive control in moving boundary problems for the wave equation, by rewriting the problem into an LTI form with time-varying input delay and unknown parameters to which adaptive backstepping designs can be applied. See also Section 5.2 for some applications of these designs.
@@ 4.7 Fredholm operators and non-strict feedback plants
As mentioned in the primer, in Section 2, while the original backstepping method was developed using Volterra integral transformations, see (8), in some cases this does not give enough degrees of freedom to find an adequate target system and well-posed kernel equation for the control design. The first specific example where a different type of transform was looked for is likely [247], to assign an arbitrary decay rate for an Euler-Bernoulli beam equation. This early example builds a control law based on a previous design for the Schrödinger equation, which uses a standard Volterra transformation, and then adapts it to obtain stabilization to zero of the Euler-Bernoulli beam equation instead of to a constant. Unlike later backstepping designs, this design depends on the system dynamics when defining the properties of the transforms, these are not seen as operators mapping from generic function spaces (e.g. to ), but relating solutions to two specific differential equations. This is reflected in the transformation structure, where the time derivative of the PDE state appears alongside the value of the state itself. In this particular case, the Fredholm transformation linking the two sets of solutions is found after imposing a particular form of the control action in order to obtain explicit expressions for the closed-loop system. The next examples in the literature are [45], with a wave PDE in the actuation path of an ODE, and [47], in the case of distributed input delays in an ODE. In these examples, the direct backstepping transformation is written in a more general form. Nevertheless, this transformation, due to the particular structure of the system under consideration, has good invertibility properties. In particular, the inverse transform only contains Volterra integrals, which seems to indicate that the transformation is not a generic Fredholm transform. Another particular structure, this time concerning parabolic equations with integral terms, was presented in [150] for a spatially non-causal reaction-diffusion equation. In this case, the particular form of the integral kernel in the non-local term is exploited to obtain an explicit solution for the Fredholm transformation and its invertibility. After these first examples, two different early approaches emerged to tackle more general Fredholm transformations, a contraction approach in [68] and [69] that required little structure from the coefficients yet only worked for “small” coefficients; and a more technically involved approach tailored to self-adjoint and skew-adjoint operators in the case of a Korteweg-de Vries equation in [91] and a Kuramoto-Sivashinsky equation in [92], respectively. In these last two papers, the existence and invertibility of the transform were related to controllability conditions. Many recent results follow more closely a later approach, presented in [88], which linked the existence and invertibility of the Fredholm transform to a controllability condition in the case of first-order differential equations, yet required less specific conditions on the operators studied.
@@ 4.8 Sampled-data and Event-triggered systems
As discussed in section 4.1, one of the strengths of the backstepping method has been its capacity to simplify convoluted structures in PDE systems and allow for an equivalent stability analysis based on a simpler target system. This advantage has been exploited not only in adaptive and parameter estimation methods but also in sampled-data and Event-triggered stabilization and estimation schemes, which generally exploit the easier structure of the target system in order to construct Lyapunov functions for the closed-loop system (or to estimate the rate of decay of some norm of the original system).
@@
Event-based control results include linear hyperbolic systems in [131], as well as an observer-based version in [129]. Another related system can be found in [130], where traffic flow on cascaded roads is considered under event-triggered output-feedback. As for parabolic PDEs, existing results include [132] for reaction-diffusion PDEs with constant coefficients, as well as [230] for an observer-based control and [231] considering also sampled-data. For interconnected PDE-ODE systems, some results exist, such as [205] in the hyperbolic case, and [289] in the parabolic case.
@@ 5.1 Traffic Control
Freeway traffic modeling and control have been intensively investigated over the past decades. Backstepping has proven to be an effective method for designing boundary controllers that address stop-and-go oscillations in congested traffic, often modeled by the nonlinear Aw-Rascle-Zhang (ARZ) model [305, 303]. This model, consisting of second-order hyperbolic PDEs, accurately captures the complex dynamics of traffic density and velocity. Early backstepping designs focused on stabilizing traffic flow on a single freeway segment, considering both upstream and downstream control problems using ramp metering [305]. They employed backstepping transformations to map the original system to a target system with desirable stability properties, yielding full-state feedback control laws and collocated/anti-collocated observers. These designs achieved exponential stability and finite-time convergence to equilibrium, validated through simulations. Later works expanded the methodology to address the more challenging problem of simultaneous downstream and upstream stabilization on cascaded freeway segments connected by a junction, considering varying road conditions [303]. This scenario requires a more sophisticated control approach due to the interconnected nature of the system and distinct traffic scenarios on different segments. Backstepping-based output-feedback controllers were designed to suppress stop-and-go oscillations, leveraging boundary measurements of flow rate and velocity at the junction or the outlet of the segment. These designs ensured robust stabilization of the under-actuated network of two hyperbolic PDE systems, effectively mitigating congestion on cascaded roads. In [148, 149], these works further extended the application of backstepping to the control of multi-class traffic flow, incorporating the influence of disturbances and bottlenecks, they designed an optimal observer-based output feedback controller for a heterogeneous traffic flow model with disturbances at the inlet of a considered road segment. Leveraging the backstepping approach, the proposed controller guaranteed the integral input-to-state stability of the closed-loop system and effectively addressed traffic breakdown caused by a combination of high traffic demand, bottlenecks, and driver disturbances. The controller design involved constructing an observer that estimates the states of the linearized ARZ traffic flow model using only boundary measurements. In [147], the backstepping method was employed to design an observer for a heterogeneous quasilinear traffic flow system, explicitly addressing the impact of disturbances at the inlet of the road section. This work demonstrated the applicability of backstepping for state observation in the presence of perturbations, ensuring accurate estimation of the traffic state despite uncertainties and nonlinearities. Backstepping has also been applied to address the control of traffic flow with multiple vehicle classes, each characterized by its own size and driver behavior [70]. This requires a multi-class traffic model, typically represented by a set of coupled hyperbolic PDEs. By designing an output-feedback controller based on backstepping and an anti-collocated observer, researchers achieved finite-time damping of stop-and-go waves by measuring the velocities and densities of different vehicle classes at the inlet of the considered track section. In another work, a backstepping observer was developed and validated with real freeway data to estimate congested freeway traffic states based on the ARZ model [304]. This observer design only requires boundary measurements of flow and velocity, making it suitable for practical applications with limited sensor coverage. Finally, recent contributions used the backstepping approach to address the problem of mean-square stabilization of mixed-autonomy traffic PDE systems [313].